OpenStudy (anonymous):

Can someone please explain "standard form" for me? I never learned it even though I was supposed to- my teacher told me: Ax+By=C. What does it solve? What am I supposed to do with it? Explain anything you can please xx

OpenStudy (anonymous):

It's another way of writing the equation of a line. It's not particularly useful, but if your teacher tells you to write the equation of a line in standard form, just rearrange terms until you get it in standard form.

OpenStudy (anonymous):

Can you show me an example? (step by step) *It's you again! xD haha*

OpenStudy (anonymous):

So, an example: y=2x+3. Subtracting 2x gives you y-2x=3. That is in the form of Ax+By=C.

OpenStudy (anonymous):

Can be written as -2x+y=3. Same thing.

OpenStudy (anonymous):

Soo it's just like y=mx+b (a way to write a linear equation). Is there a way I can graph through standard form? Is there a certain use to it?

OpenStudy (anonymous):

Actually, no. I was wrong. Supposedly the coefficient of the x-term must be positive. SO in my last example, make x positive by multiplying each term by (-1). Thus, (-1)(-2x) is now 2x, (-1)(y) is now -y, and (-1)(3) is -3. So, together: 2x-y=-3 is the right answer.

OpenStudy (anonymous):

Why does the x have to be positive?

OpenStudy (anonymous):

No, no way to use it. It's stupid. Just turn equations into that when necessary. I have no idea. Just know how to get that form, and do it if the teacher asks you to.

OpenStudy (anonymous):

So the x always has to be positive, the others don't matter?

OpenStudy (anonymous):

Also, my teacher said that A, B and C always have to be integers (whole numbers)? What's the deal with that?

OpenStudy (anonymous):

Yeah. Also, fractions are not allowed in this form. If you hae a fraction, just multiply each term by the reciprocal. That's what integers are- whole numbers (any number that's not a fraction). So if there's (1/2)(x), multiply everything by 2 (the reciprocal of 1/2). This makes x be (1)(x), and 1 is a whole number.

OpenStudy (anonymous):

An example: http://www.algebra-class.com/image-files/standard-form2.gif.

OpenStudy (anonymous):

So the x always has to be positive, the others don't matter?

OpenStudy (anonymous):

Whoops. http://www.algebra-class.com/image-files/standard-form2.gif

OpenStudy (anonymous):

Sorry didn't mean to type that.

OpenStudy (anonymous):

Okay so basically all I have to know is: reorder it to Ax+By=C, NO fractions, and x must be positive? And that this has no use?

OpenStudy (anonymous):

Other than to make the equation look pretty? No. And yes, that's all you have to know.

OpenStudy (anonymous):

They say it's useful when graphing, but slope-intercept works just as well, in my opinion.

OpenStudy (anonymous):

How is it useful when graphing?

OpenStudy (anonymous):

Okay so basically all I have to know is: reorder it to Ax+By=C, NO fractions, and x must be positive? And that this has no use?

OpenStudy (anonymous):

It isn't. Use slope-intercept form almost every time (except when it's necessary to use point-slope form).

OpenStudy (anonymous):

Ugh sorry my computer does that.

OpenStudy (anonymous):

Thanks(: http://openstudy.com/study#/updates/522d0e12e4b0fbf34d633970

OpenStudy (anonymous):

Can you see if I answered that right?